OpenStudy (anonymous):

I need help with finding out what x= to the following problem 1/3 log(x) - log (10)=1 x= ?

OpenStudy (anonymous):

\[\large a.\log(x) = \log(x)^a\]

OpenStudy (across):

Hint:\[\ \frac{1}{3}\log(x)-\log(10)=1\\ \log(x^\frac{1}{3})-\log(10)=1 \]

OpenStudy (anonymous):

Then use this formula: \[\large \log(x) - \log(y) = \log(\frac{x}{y})\]

OpenStudy (anonymous):

On right hand side you can write 1 as: \[\large \color{green}{\log(0)}\]

OpenStudy (anonymous):

I do not understand what x=

OpenStudy (rsadhvika):

another solution : \(log x -3 log 10 = 3\) \(log x -log 10^3 = 3\) \(log x/10^3 = log 10^3\) \(x = 10^6\)

OpenStudy (anonymous):

thank you

OpenStudy (rsadhvika):

its correct. but my bad!! you can solve it even simpler way

OpenStudy (rsadhvika):

log 10 = 1

OpenStudy (rsadhvika):

could you try this way and confirm my answer ?